### v2

# Dynamic thermalization on noisy quantum hardware

Abstract
Relaxation after a global quench is a natural way to probe thermalization in closed quantum systems. When a system relaxes after the quench, thermal observables emerge in the absence of constraints, provided long-time averaging or a large system. We demonstrate a thermalization mechanism based on averaging the observables over realizations of a global quench protocol that does not rely on a system’s size or long-time evolution. The quench abruptly establishes all-to-all couplings of random strength in a few-body system and initializes the dynamics. Shortly after the quench, the observables averaged over realizations of random couplings become stationary. The average occupation probabilities of many-body energy states equilibrate toward the Gibbs distribution with a finite positive or negative absolute temperature that depends on the initial state’s energy, with the negative temperatures occurring due to the confined spectrum of the system. Running an experiment on an IBM Quantum computer (IBMQ) for four qubits, we report the utility of the digital quantum computer for predicting thermal observables and their fluctuations for positive or negative absolute temperatures. Implementing thermalization on IBMQ, this result facilitates probing the dynamical emergence of thermal equilibrium and, consequently, equilibrium properties of matter at finite temperatures on noisy intermediate-scale quantum hardware.
The notions of temperature and statistics in isolated quantum systems are based on the mean and variance of a physical observable over repeated experiments [1, 2]. Microscopically, the emergence of statistical mechanics from the quantum description of a closed system is understood via thermalization. Thermalization implies that the quantum mechanical expectation value of the observable dynamically relaxes to the thermal one [3, 4, 5, 6, 7]. Achieving thermalization requires the system to equilibrate: the expectation value of an observable $\cal O$ is stationary,
$\langle{\cal O}\rangle_{\rm eq}=\overline{\langle{\cal O}\rangle}={\rm Tr}\,\overline{\rho}\,{\cal O},$ (1)
where $\overline{\rho}$ is the time-averaged diagonal density matrix and the trace is taken over the entire system. In turn, the equilibrium expectation value (1) is thermal if it agrees with the one evaluated with the thermal density matrix $\rho_{\rm th}$ describing a Gibbs state,
${\rm Tr}\,\overline{\rho}\,{\cal O}={\rm Tr}\,\rho_{\rm th}{\cal O}={\cal O}_{\rm th}.$ (2)
Relaxation after a quench is one way thermal observables occur in a closed quantum system. The quench induces an interaction between the system’s subparts and initiates the dynamics. As a result, the system equilibrates (1) and thermalizes (2) unless it possesses extra constraints that, e.g. result in integrability or lead to many-body localization [8].
In the present research, we demonstrate an alternative microscopic mechanism for the occurrence of stationary observables and the following thermalization in the course of post-quench dynamics. For an isolated system of several non-degenerate qubits, we consider a quench protocol that turns on an all-to-all interaction with coupling constants sampled from a complex Gaussian distribution with zero mean and finite variance. The protocol is repeated multiple times with different sets of couplings. We find the equilibrium value of an observable dynamically arising from averaging over the realizations of the quench protocol:
$\langle{\cal O}\rangle_{\rm eq}=\langle\langle{\cal O}(t>t_{\rm eq})\rangle\rangle.$ (3)
The double brackets denote the quantum mechanical mean averaged over realizations, and $t_{\rm eq}$ is the time after which the observable becomes stationary. The observable equilibrates to the thermal value,
$\langle\langle{\cal O}(t>t_{\rm eq})\rangle\rangle={\cal O}_{\rm th}(T(\varepsilon_{\rm in})),$ (4)
where the temperature $T$ appears due to the quenched interaction and depends on the energy of the initial state $\varepsilon_{\rm in}$. The system thermalizes to a positive or negative temperature state as it has a bounded spectrum and no kinetic energy [9]. Below, we demonstrate this approach leads to thermal observables for the pre-quenched system without an external thermal bath.
We experimentally examine our thermalization mechanism (3,4) in a simulation on an IBM Quantum computer (IBMQ) for a system of four qubits. Upon equilibration, we read out the occupation probabilities of the many-body energy states of the system described by the pre-quench Hamiltonian. For the system’s initial state given by the ground state of the qubits, the average occupation probabilities relax to the Gibbs distribution. Starting from the highest excited state leads to the inversely populated state ($T<0$). As a result, both the observables and their fluctuations evaluated from the noisy quantum computer’s error-mitigated output display thermal behavior with positive or negative absolute temperatures.
Being implemented on noisy intermediate-scale quantum (NISQ) hardware, our approach allows probing thermalization in finite-size quantum systems in a highly controlled environment and may be utilized among nature-inspired quantum algorithms for thermal state preparation [10, 11]. Considerable experimental effort has been spent on sampling observables agreeing with the Gibbs state prediction on a digital (gate-based) quantum computer. Some recent experiments are based on a quantum imaginary time evolution algorithm [12, 13] or a variational approach with thermofield double states [14, 15], while another incorporates a thermal bath using ancilla qubits [11]. These experiments achieve thermalization for relatively small systems of $2$ to $4$ qubits. One concludes thermalization from the agreement of the observables measured at the end of the simulation with the Gibbs state prediction. Here, we also address thermalization in a small system on a digital quantum computer. We rely on the post-quench evolution and averaging the observables over realizations of the randomized auxiliary interaction. Our protocol is inherently dynamical, enabling studying the formation of thermal observables in real time. Furthermore, it relies on classical averaging of quantum observables corresponding to an individual realization of the quench protocol, unloading precious resources of the quantum simulator. Advantageously, it does not require long-time evolution or increasing the system size due to auxiliary qubits aiming to simulate a thermal bath.
[…]
The model.— Quenching a closed system is a natural way of probing its equilibration dynamics and subsequent thermalization [16]. For example, thermalization may occur via a non-linear classical mode and the development of turbulence, leading to eventual energy equipartition [17], or due to switching on a highly non-local interaction in a quantum system resulting in its equilibration to an infinite temperature state [18, 19]. Taking this dynamical perspective, we organize our quench protocol as follows.
We prepare $N$ non-degenerate qubits in an eigenstate of $H_{0}=\sum_{j=1}^{N}(\omega_{j}/2)\,\sigma^{y}_{j}$: $H_{0}|\psi_{\rm in}\rangle=\varepsilon_{\rm in}|\psi_{\rm in}\rangle$. Here $\sigma_{j}^{y}$ is the second Pauli matrix acting on the $j$th qubit, $\omega_{j}$ is the $j$th qubit’s frequency, $|\psi_{\rm in}\rangle$ is the initial state of the system, and $\varepsilon_{\rm in}$ is the corresponding eigenenergy. At time $t=0$, we abruptly turn on a global quench that couples the qubits all-to-all and initiates the dynamics. The initial state $|\psi_{\rm in}\rangle$ starts evolving with the Hamiltonian
$H=H_{0}+{\cal V},$ (5)
where
${\cal V}=-\sum_{j_{1}>j_{2}}^{N}J_{j_{1}j_{2}}\sigma_{j_{1}}^{+}\prod^{j_{1}-1}_{k=j_{2}+1}\!\sigma_{k}^{z}\,\sigma_{j_{2}}^{-}+h.c.$ (6)
describes the interaction between the qubits with the coupling constants $J_{j_{1}j_{2}}$ drawn from the Gaussian unitary ensemble with zero mean $\langle J_{j_{1}j_{2}}\rangle=0$ and finite variance $\langle|J_{j_{1}j_{2}}|^{2}\rangle=J^{2}/N$. Accordingly, one may understand the interaction term (6) as the Hamiltonian of random non-interacting fermions [20] projected onto the qubits’ basis via the Jordan-Wigner transform. While both $H_{0}$ and ${\cal V}$ are integrable their combination is not. Indeed, when ${\cal V}$ is combined with $H_{0}$ in the full Hamiltonian (5), the latter is chaotic displaying Wigner-Dyson level statistics [21] as discussed in more detail in Appendix A. As such, the Hamiltonian (5) describes an interacting theory from fermions’ or qubits’ perspective giving rise to the quantum many-body dynamics we address below.
We repeat the time evolution of the initial state $|\psi_{\rm in}\rangle$ with the Hamiltonian (5) for different sets of random couplings and compute quantum mechanical mean values averaged over realizations of the quench protocol, aiming to check the equilibration (3) and thermalization (4) conditions.
Once the interaction ${\cal V}$ is on, the evolving state $|\psi(t)\rangle=e^{-it(H_{0}+{\cal V})}|\psi_{\rm in}\rangle$ is no longer an eigenstate of $H_{0}$ whose expectation value $\langle H_{0}(t)\rangle=\langle\psi(t)|H_{0}|\psi(t)\rangle$ becomes dynamic. To study the dynamics, we write $H_{0}=\sum_{l}\varepsilon_{l}|l\rangle\langle l|$ in the basis of its $L=2^{N}$ many-body energy eigenstates labeled by the index $l$, and average $\langle H_{0}(t)\rangle$ over the protocol’s realizations. Thereby, we have
$\langle\langle H_{0}(t)\rangle\rangle=\sum_{l=0}^{L-1}\varepsilon_{l}\langle\langle n_{l}(t)\rangle\rangle,$ (7)
where
$\langle\langle n_{l}(t)\rangle\rangle=\langle\langle l|e^{-it(H_{0}+{\cal V})}|\psi_{\rm in}\rangle\langle\psi_{\rm in}|e^{it(H_{0}+{\cal V})}|l\rangle\rangle,$ (8)
is the average occupation probability of the many-body state with energy $\varepsilon_{l}$ for $t>0$. Throughout the evolution, the total energy of the system is conserved on average, $\langle\langle H(t)\rangle\rangle=\langle\langle\psi_{\rm in}|H|\psi_{\rm in}\rangle\rangle=\varepsilon_{\rm in}$, due to the zero mean of the Gaussian in Eq. (6).
Exact dynamics.— We choose $N\!=\!4$ and run the complete protocol using exact diagonalization for different initial states $|\psi_{\rm in}\rangle\!=\!|\!\downarrow\downarrow\downarrow\downarrow\rangle,|\!\uparrow\downarrow\downarrow\downarrow\rangle,\ldots,|\!\uparrow\uparrow\uparrow\uparrow\rangle$, all having distinct initial energies, where the down arrow denotes a qubit in the ground state of $\sigma^{y}_{j}$ and the up arrow marks the excited state of a qubit.
As shown in Fig. 1a, the average expectation value (7) becomes stationary at $t_{\rm eq}\approx 2.5/J$ up to small finite-size oscillations. On the contrary, the expectation value for each realization of random couplings persistently oscillates at high amplitude. The equilibration of the mean (7) appears along with the building up of the diagonal entropy, $S(t)=-\sum_{l}\langle\langle n_{l}(t)\rangle\rangle\ln\langle\langle n_{l}(t)\rangle\rangle$, which agrees with the properties of the thermodynamic entropy [22] and approaches an average stationary value $S(t>t_{\rm eq})$ in Fig.  1b. Alongside, in Figs. 1a, we note that the equilibrated observables depend on the choice of the initial state. To rule out a possibility that our simulation has reached a quasi-equilibrium steady state, such as prethermalization [23, 24], rather than fully equilibrated, we explore the dynamics of the observables at times far exceeding $t_{\rm eq}$. Even though prethermalization is often associated with a weak breaking of integrability, it was recently observed in a trapped ion system with no conserved charges [25]. In this case, achieving eventual thermalization and losing the information of the initial state requires evolution up to a timescale of the inverse minimal level spacing [26]. In Fig. 1c, we demonstrate that the average expectation value (7) persists at times far exceeding the inverse minimal many-body level spacing of $H$, indicating complete equilibration, yet remains dependent on the initial state, which reflects the average total energy’s conservation.
Having established equilibration (3), we explore the relaxation process further. Since the Hamiltonian $H=H_{0}+{\cal V}$ is chaotic, it is natural to expect that the system equilibrates and thermalizes if it is large enough [6]. For a small system of $N=4$ qubits, a dynamic observable averaged over an ensemble of realizations equilibrates after passing an interaction-dependent timescale $t_{\rm eq}$, even though an observable computed with an individual realization of the quench protocol may severely oscillate. Generally, one can present the evolved state of the system as $|\psi(t)\rangle=e^{-i(H_{0}+{\cal V})t}|\psi_{\rm in}\rangle=\sum_{l}c_{l}(t)|l\rangle$, where we specifically choose the basis of eigenstates of the initial Hamiltonian $H_{0}$. Accordingly, the coefficients of this representation $c_{l}(t)=\langle l|\psi(t)\rangle$ are related to the probability of occupying the $l$-th state at time $t$: $\langle n_{l}(t)\rangle=|c_{l}(t)|^{2}$. For $t>0$, these occupation probabilities begin to deviate from their initial values and become dynamically redistributed, evolving with the chaotic Hamiltonian (5). If the system equilibrates, the probabilities to occupy the $l$-th state of $H_{0}$ should become stationary, describing an equilibrium probability distribution. Here we check if the average occupation probabilities become thermal.
Generically, describing the equilibrium state of a quantum many-body system requires a generalized Gibbs ensemble to account for the conserved quantities. The generalized Gibbs ensemble reduces to the conventional Gibbs ensemble if the conserved quantities are limited to total energy, particle number, and momentum [27]. Since our quench protocol respects only the total energy conservation on average, we compare the averaged occupation probabilities upon equilibration with those predicted by the usual Gibbs distribution:
$n_{l}^{\rm th}(T)=\frac{e^{-\varepsilon_{l}/T}}{Z},\quad Z(T)=\sum\limits_{l=0}^{L-1}e^{-\varepsilon_{l}/T}.$ (9)
By doing that, we aim to achieve thermal occupations for the initial system, assuming that the emergent temperature $T$ absorbs the effects of the interaction. At $t=10/J>t_{\rm eq}$, we observe that while the occupation probabilities $\langle n_{l}(t)\rangle$ stay irregular for each individual realization of the quench protocol, the average occupation probabilities (8) are well-approximated with the occupation probabilities $n_{l}^{\rm th}(T)$. To illustrate this agreement, we fit the average occupation probabilities (8) with the thermal ones (9) obeying the Gibbs distribution with the temperature $T$ being a fitting parameter. We compare $\langle\langle n_{l}(t)\rangle\rangle$ with the thermal fit $n_{l}^{\rm th}(T)$ in Fig. 1d.
As expected upon thermalization, the occupations depend on a single parameter – temperature, which, in our case, is set by the energy of the initial state. One may adjust the temperature’s absolute value by tuning the characteristic interaction strength, providing additional control over the thermalization process. The temperature retains the memory of the total energy persisting as a constraint throughout the relaxation, which makes the equilibrium observables initial-state-dependent. Notably, the choice of the initial state also defines the temperature’s sign, giving rise to negative absolute temperatures. Plotting equilibrated mean occupation probabilities for the protocol started from the ground state (Fig. 1d left) and the highest excited state (Fig. 1d right) of the qubits, we find a population inversion in the second case, which implies $T<0$. While temperature is usually restricted to be non-negative for an unbounded energy spectrum to secure convergence of the partition function, negative absolute temperatures have been experimentally observed in systems with bounded spectra [28, 29, 30, 31, 32, 33, 34, 35, 36], as in our simulation.
To justify the quantitative agreement between the mean occupation probabilities and the Gibbs distribution, we consider the Kullback-Leibler (KL) divergence [37]. The KL divergence serves as a measure of distance between the two probability distributions. It is non-negative by definition and equals zero only if the two probability distributions coincide. Since the average observables still possess small finite-size oscillations, as seen in Fig. 1a, to find the equilibrium temperature $\overline{T}$, we additionally average the mean occupations over a fixed time interval in the equilibrium regime (from $t=2.5/J$ to $t=20/J$) and approximate the result with the thermal distribution (9) with $T=\overline{T}$. Then, we compare $n_{\rm th}(\overline{T})$ with the dynamic occupations (8) via the KL divergence [38, 39]:
$D(\langle\langle n_{l}(t)\rangle\rangle||n_{l}^{\rm th}(\overline{T}))\!=\!-S(t)\!+\!\frac{\langle\langle H_{0}(t)\rangle\rangle\!-\!F(\overline{T})}{\overline{T}},$ (10)
where $F(\overline{T})\!=\!-\overline{T}\ln Z(\overline{T})$ is the thermodynamic free energy. For several initial states, we determine $\overline{T}$ and plot the KL divergence (10) as a function of time after the quench in Fig. 1e. The KL divergence decays for $t>t_{\rm eq}$ to the value $\sim 10^{-2}$ which quantifies the accuracy of the thermal occupation probabilities.
[…]
Figure 2: Thermalization dynamics of a four-qubit transverse Ising model with a periodic boundary condition in a paramagnetic phase after quenching ${\cal V}$. We use $\omega/g=4$ and $g=0.05J$. a) Expectation value of $H_{I}$ as a function of time after the quench. The gray curves show individual realizations of random couplings. The blue curve is the expectation value of $H_{I}$ averaged over $200$ realizations of random couplings. b) Probabilities to occupy eigenstates of $H_{I}$ read out upon equilibration. Alike in Fig. 1d, the blue squares show the average occupation probabilities (8) and red dots are the thermal fit (9). The spectrum of $H_{I}$ contains two pairs of degenerate states at $\varepsilon_{l}=\pm 0.4$ and a $4$-degenerate state at $\varepsilon_{l}=0$. c) Kullback-Leibler divergence between the average occupation probabilities and thermal occupations with the equilibrium temperature $\overline{T}\approx 0.87J$.
So far, we have achieved thermal occupations of the many-body energy states for the system described by the Hamiltonian $H_{0}$ of four non-interacting qubits, starting the protocol from an eigenstate of $H_{0}$. This setup can not evolve to a thermal equilibrium without the interaction potential ${\cal V}$ that breaks integrability and leads to eventual thermalization on average. To highlight the generality of our thermalization protocol, we choose an $N=4$ transverse Ising model with a periodic boundary condition, $H_{I}=-\omega\sum_{j=1}^{N}\sigma_{j}^{x}-g\sum_{j=1}^{N-1}\sigma_{j}^{z}\sigma_{j+1}^{z}-g\sigma_{N}^{z}\sigma_{1}^{z}$, as an initial Hamiltonian. This system is also integrable, yet the induced interaction (6) again leads to average thermal occupations of the eigenstates of $H_{I}$ with eigenenergies $\varepsilon_{l}$, as shown in Fig. 2. We tune the initial Hamiltonian in to a paramagnetic phase ($\omega=0.2J$, $g=0.05J$) and begin the evolution with a product state, $|\psi_{\rm in}\rangle=(1/\sqrt{2}\begin{pmatrix}1&1\end{pmatrix}^{\!\rm T})^{\otimes N}$, which is no longer an exact eigenstate of $H_{I}$: $\varepsilon_{\rm in}\neq\varepsilon_{l}$. Running $200$ realizations of random couplings, we observe in Fig. 2a that an average expectation value of $H_{I}$ equilibrates at $t\sim{\cal O}(1)/J$, where $J$ is the largest energy scale in the problem. We plot the occupation probabilities for $t=25/J$ in the equilibrium regime in Fig. 2b and find that mean occupation probabilities are well approximated by the Gibbs distribution. Note that the spectrum of $H_{I}$ has multiple degeneracies. Finally, in Fig. 2c, we show the Kullback-Leibler divergence (10) between the mean dynamic occupations and Gibbs distribution with an equilibrium temperature $\overline{T}$ determined from fitting the dynamic data averaged over an interval from $t=10/J$ to $t=50/J$ by thermal ones (9). These results are qualitatively similar to our previous analysis, namely, Figs. 1a,d,e.
Even for a small system of four qubits, thermalization to a finite positive or negative temperature equilibrium occurs from the post-quench dynamics of a quantum system within the time scale $\sim{\cal O}(1)/J$. Suitability for a small system, controllable equilibration timescale, and a lack of a thermostat make our thermalization mechanism possible to test experimentally on NISQ computers like IBMQ. Indeed, as we demonstrate below, executing the evolution of the interacting four-qubit system, described by the Hamiltonian (5), is viable on IBMQ with a moderate number of gates that often is the main limiting factor in studying many-body physics on NISQ devices. Achieving thermal observables relies on averaging the expectation values, which can be done classically upon readout from every realization of the quench protocol, each running independently on a quantum computer.
Quantum simulation.— We implement our dynamic thermalization protocol on an IBM Quantum computer (IBMQ) named ibm_hanoi for $100$ realizations of the quench protocol, each corresponding to a quantum circuit. Encoding quantum dynamics into a quantum circuit implies applying a sequence of single- and two-qubit gates that approximate a given realization of the evolution operator $U(t)=e^{-it(H_{0}+{\cal V})}$ to the initial state $|\psi_{\rm in}\rangle$. Executing a quantum circuit on IBMQ is noise-susceptible, mostly due to the imperfect application of two-qubit gates, namely, CNOT gates that perform entangling operations between qubits in our simulation. Hence, our strategy to mitigate the error accumulated due to noise is: first, decreasing the number of CNOT gates required to perform the simulation as much as possible; second, simplifying the noise’s structure from the remaining CNOT gates and accounting for it in the simulation results.
Figure 3: Quantum simulation results for evolution up to $t=10/J$ on the IBM Quantum computer. a) Occupation probabilities averaged over $100$ realizations of the quench protocol with the initial state $|\psi_{\rm in}\rangle=|\!\downarrow\downarrow\downarrow\downarrow\rangle$: the gray crosses display the result of the exact dynamics (ED); the dash-dotted gray line is the thermal fit (9) of the ED data; the blue circles show the outcome of the quantum simulation (QC) on IBMQ after error mitigation; the solid blue curve is the fit of the QC data with the model for noisy occupations (11); the red diamonds and the dashed red curve show the same as QC but with tripled noise amplification (QC3). b) The Gibbs distribution for occupation probabilities (9) with the temperatures determined from fitting the ED and quantum simulation data in the panel a. The error bars for QC and QC3 data sets in panels a and b are within the size of the designated symbols. ZNE refers to the temperature value obtained with zero-noise extrapolation for $\beta=1/T$ shown in the panel c. Panels d, e, f are analogous to a, b, c but for the protocol starting with $|\psi_{\rm in}\rangle=|\!\uparrow\uparrow\uparrow\uparrow\rangle$. mQC3 is the mirror image of the QC3 value of the inverse temperature $\beta=1/T$ relative to the QC value of $\beta$, necessary for mirrored zero-noise extrapolation (mZNE) illustrated with the dashed magenta line in the panel f. The frequencies of the qubits are the same as in Fig. 1. The plots for occupation probabilities are in the log scale along the y-axis.
To fulfill the first task, we adapt variational circuit recompilation [40, 41] to approximate the evolution operator $U(t)$ by a parametrized quantum circuit with $99.8\%$ accuracy, optimizing the tensor network decomposition of $U(t)$, before plugging it into the quantum computer (see Appendix B). Using this method, we reduce the number of CNOT gates required for the simulation to $N_{\rm CNOT}=60$, discarding the need for Trotterization, which enables one to reach arbitrary times.
Figure 4: Exact vs Quantum dynamics for the expectation value of $H_{0}$ and its fluctuations. a) Average expectation value of $H_{0}$ for $|\psi_{\rm in}\rangle=|\!\downarrow\downarrow\downarrow\downarrow\rangle$. ED denotes the exact dynamics, EDβ approximates the ED curve with the thermal fit, and ZNE refers to the observables computed with thermal occupations obtained from the quantum computer in the zero-noise limit. b) Fluctuations of $H_{0}$, $\sqrt{\langle\langle H_{0}^{2}(t)\rangle\rangle-\langle\langle H_{0}(t)\rangle\rangle^{2}}$, for $|\psi_{\rm in}\rangle=|\!\downarrow\downarrow\downarrow\downarrow\rangle$. Panels c and d show the same as in panels a and b but for the different initial state, $|\psi_{\rm in}\rangle=|\!\uparrow\uparrow\uparrow\uparrow\rangle$. mZNE designates the observables computed with the output of the mirrored ZNE approach we used to establish the zero-noise limit for states with negative temperatures.
We address the second task using an extension of the randomized compiling (RC) approach [42] converting the noise, not only on qubits to which a CNOT gate is applied but also on the neighboring qubits, to an incoherent noise channel (see Appendix C). The incoherent noise contains local and global components, where the latter acts uniformly on every qubit in the circuit. To account for the global incoherent noise, we approximate the noisy quantum computer’s output for average occupation probabilities (8) upon equilibration as a weighted mixture of the average occupations unaffected by the global incoherent noise and a uniform distribution describing a maximally mixed (infinite temperature) state. Aiming to see if the observables evaluated on the quantum computer display thermal behavior, we chose the average occupations unaffected by the global incoherent noise in the form of the finite temperature Gibbs distribution (9) so that the noisy average occupation probabilities read
$\langle\langle n_{l}(t>t_{\rm eq})\rangle\rangle=f^{N_{\rm CNOT}}n_{l}^{\rm th}(T)+\frac{1-f^{N_{\rm CNOT}}}{2^{N}}.$ (11)
The consistency of this choice settles by fitting the noisy quantum computer’s data to the model (11) where the temperature $T$ and the circuit’s fidelity $f^{N_{\rm CNOT}}$ play a role of the fitting parameters.
[…]
The temperature obtained from the quantum simulation is higher than the temperature from the thermal prediction of the exact dynamics, as we observe from the occupations’ slopes in Fig. 3b. We attribute this heat-up to the part of the hardware noise so far unattended. Indeed, earlier, we have discussed the effect of the global incoherent noise, yet the local incoherent noise may be present in the system. We account for the remaining noise using the zero-noise extrapolation (ZNE) method [43, 44, 45, 46, 47, 48], which involves repeating the experiment with artificially amplified noise and extrapolating the measurement results to the noiseless limit. For incoherent noise, a linear change in noise amplification generally leads to exponential change in an observable [45]. From this perspective, the inverse temperature $\beta=1/T$ in the exponent of thermal occupations in Eq. (11) may absorb the impact of the remaining noise. Elaborating on this observation, we repeat the experiment with the tripled noise by tripling every CNOT gate in the circuit. The average occupation probabilities in the quantum simulation with tripled noise amplification (QC3) remain well within the incoherent noise approximation, (11) as shown in Fig. 3a. In turn, the temperature determined from the QC3 data (see Fig. 3b) increases further compared to the QC result, highlighting how the hardware noise on NISQ devices can lead to the temperature rise in the simulated observables. In Fig. 3c, we extrapolate the inverse temperature to the noiseless case using a linear fit based on two noise amplification values (QC, QC3). As the inverse temperature appears in the exponent of the noisy occupations (11), our method effectively performs an exponential extrapolation on the observables after clearing the effect of the global incoherent noise from the quantum data. As a result, the value of $\beta_{\rm ZNE}=(1.29\pm 0.12)/J$ determined from ZNE lies within the error bar with $\beta_{\rm ED}=(1.21\pm 0.04)/J$ found from fitting the occupations computed in the exact dynamics with the Gibbs distribution (9). The canonical prediction for $\beta_{c}=1/T_{c}=1.20/J$ based on energy conservation well agrees with exact dynamics and experimental value of the inverse temperature upon error mitigation.
The state evolved from the highest energy state, $|\psi_{\rm in}\rangle=|\!\uparrow\uparrow\uparrow\uparrow\rangle$, equilibrates to the inversely populated state in quantum simulation as it does for the exact dynamics with the QC results well-described by the noisy occupation probabilities (11) with $T<0$, as shown in Figs. 3d,e. Tripling the noise on the quantum computer leads to the system’s equilibration to a state with $T>0$. Indeed, negative temperature states are unstable when interacting with an environment [9], which makes them particularly challenging to prepare. Since negative absolute temperatures are hotter than $T=+\infty$, with $T=0^{-}$ being the hottest [9, 31], from a thermodynamic standpoint, the environment (considered at $T>0$) cools the system at $T<0$ to a positive temperature. Thinking of the hardware noise as a consequence of a quantum computer being an open system, one expects the inverse populations to be unstable against noise, as we observe for the QC3 occupation probabilities in Figs. 3d,e. However, in our case, the noise plays a substantial role in the dynamic formation of the system’s equilibrium state, being determined by the hardware and the circuit layout rather than describing an independent thermal entity interacting with the system equilibrated to a negative absolute temperature. Hence, it is hard to predict a threshold noise preserving $T<0$ equilibrium. Therefore, doing zero-noise extrapolation to the noiseless negative temperature case requires caution. Upon tripling the noise, the temperature crosses infinity ($\beta=0$) and has a high positive value (small positive $\beta$), implying cooling the system. Thereby, ZNE for the QC and QC3 inverse temperatures predicts the value of $\beta_{\rm ZNE}=(-4.74\pm 0.37)/J$ quite distant from the ED result, $\beta_{\rm ED}=(-1.25\pm 0.04)/J$, in Fig. 3f. To extrapolate to the regime of a closed system, we take a heuristic approach and assume that the increasing noise heats the system monotonically, restricting $T$ to the $T<0$ sector and preventing relaxation to $T>0$ equilibrium. In this way, we account for $\Delta\beta=|\beta_{\rm QC}-\beta_{\rm QC3}|$ and mirror the QC3 data point relative to the QC accordingly. Then we do ZNE for the mirrored QC3 (mQC3) and QC points, shown in the dashed magenta line in Fig. 3f. The value of $\beta_{\rm mZNE}=(-1.15\pm 0.41)/J$ determined with this mirrored ZNE (mZNE) method is within the error bar from the result of the exact dynamics.
Having obtained the average thermal occupation probabilities in the zero-noise limit, we check if the quantum computer predicts thermal values of the dynamic observables in the equilibrium regime. We repeat the protocol for times $t=\{2.5,5,10\}/J$ and compare the IBMQ results with the exact dynamics and the thermal approximation of ED (EDβ) for the average expectation value $\langle\langle H_{0}(t)\rangle\rangle$ and its fluctuations $\sqrt{{\rm Var}H_{0}}=\sqrt{\langle\langle H_{0}(t)^{2}\rangle\rangle-\langle\langle H_{0}(t)\rangle\rangle^{2}}$. For the quench protocol started with the ground state of $H_{0}$, $|\psi_{\rm in}\rangle=|\!\downarrow\downarrow\downarrow\downarrow\rangle$, the mean expectation value (7) computed with the thermal occupations predicted by the quantum computer agrees with the ED and EDβ results within the error bars, as seen in Fig. 4a. As for the fluctuations of $H_{0}$ shown in Fig. 4b, the quantum computer’s effective output lies right on the thermal curve, indicating that the fluctuations computed on a quantum computer are closer to the thermal equilibrium value than the prediction of the exact dynamics. For $|\psi_{\rm in}\rangle=|\!\uparrow\uparrow\uparrow\uparrow\rangle$, the quantum computer predicts negative temperatures upon equilibration. In Figs. 4c,d, we demonstrate that the IBMQ output provides for $\langle\langle H_{0}\rangle\rangle$ and $\sqrt{{\rm Var}H_{0}}$ within the error bars from both ED and EDβ predictions for $T<0$ regime.
Discussion.— We show how thermal observables can dynamically emerge in a closed quantum system as a result of averaging over repeated instances of a global quench initializing Gaussian random non-local interactions between the system’s subparts. After the quench, the system pursues evolution governed by the chaotic Hamiltonian falling into the class of Gaussian unitary ensemble of random matrices. Averaging the observables over many realizations of the quench protocol leads to their equilibration (3) and thermalization (4) in time $t_{\rm eq}\sim{\cal O}(1)/J$, where $J$ is the characteristic interaction strength. The resulting averaged probabilities of occupying eigenstates of the initial system in the equilibrium regime, $t>t_{\rm eq}$, follow the Gibbs distribution with positive or negative absolute temperatures depending on the initial state’s energy, conserved on average in the course of after-the-quench dynamics. One can additionally tune the absolute value of equilibrium temperature by adjusting the interaction strength $J$. This dynamic thermalization is particularly convenient to examine on NISQ devices such as IBMQ since it does not rely on the long-time evolution, large system size, or auxiliary qubits to emulate a thermostat but solely on the repetition of the quench protocol with different sets of random coupling constants.
Adapting the established error-mitigation techniques, we implement our thermalization protocol on the IBMQ machine using the circuit depth of $60$ CNOT gates. We obtain thermal occupations for the sixteen many-body energy states of the four-qubit system, populated according to positive or negative absolute temperatures depending on the choice of the initial state. Respectively, the observables and their fluctuations found in quantum simulation agree with the thermal predictions. This utility of the digital quantum simulator for producing thermal observables may be attributed to the notion of the quantum computer as an open system and our implementation of the thermalization protocol. First, we average the expectation values over realizations of the quench protocol aiming at mimicking the thermodynamic limit and facilitating thermalization in a relatively small system. One way to think of thermalization in a small system is to consider it as a subpart of a much bigger system in the thermodynamic limit, with the rest of the system acting as a thermal bath. A recent observation notes that averaging the evolution of a system with randomized interaction results in effective dissipative dynamics [19]. Another recent study shows that adding a disorder in a simulated system in analog quantum simulations substantially reduces the necessary size of a bath for infinite temperature thermalization [49]. In our experiment, we start with a pure state and observe dynamical relaxation to a finite temperature equilibrium for averaged observables without involving a thermal bath, with the mitigated noise seemingly falling into the same effectively dissipative channel and raising the temperature of the simulated observables. Second, we apply zero-noise extrapolation that resembles taking a dissipationless limit in an open quantum system. Genuinely, these two limits are non-commutative, and their correct order, naturally implemented in our protocol, may be essential to justify the occurrence of the fluctuation-dissipation theorem in a quantum system [50]. Alongside, in our simulation, the time for the observables to equilibrate is ${\cal O}(1)/J$, similar to the timescale of irreversible dynamics $T_{3}$ in an open quantum system, bypassing which an unconstrained system thermalizes [51].
The advantage of the considered quench protocol is its inherently dynamic nature, which allows studying the occurrence of thermalization or ergodicity breaking in quantum systems in real-time, efficiently engaging the limited resources of NISQ processors. We apply our quench protocol to study the emergence of the negative absolute temperature equilibrium on a digital quantum platform. While absolute temperatures are usually restricted to positive values due to kinetic energy, equilibrium at negative absolute temperatures is possible under certain conditions, in systems with a bounded energy spectrum. Bringing negative absolute temperature equilibrium states to reality inspired significant experimental efforts due to their non-trivial thermodynamic properties, e.g., losing entropy when absorbing energy. By now, negative temperature states were experimentally observed in a few physical systems such as nuclear spins [28, 29], cold atoms [30, 31, 34], two-dimensional quantum superfluids [32, 33], optical fibers [35], and a two-level system environment in superconducting qubits [36]. Here, we present an equilibrium state with negative absolute temperature arising from the relaxation process in a quantum system realized digitally on a quantum computer, demonstrating the utility of the existing quantum processors for current research in statistical mechanics.
Acknowledgments.— We thank Alexander Balatsky, Igor Gornyi, Archana Kamal, Alexander Mirlin, Horacio Pastawski, Anatoli Polkovnikov, and Alexander Shnirman for stimulating discussions. HP acknowledges support from the Horizon Europe program HORIZON-CL4-2021-DIGITAL-EMERGING-01-30 via the project 101070144 (EuRyQa). HP and TS acknowledge funding from the state of Baden-Württemberg through the Kompetenzzentrum Quantum Computing Project QC4BW. AIP was supported by the German Ministry of Education and Research (BMBF) within the project QSolid (FKZ: 13N16151). NVG was supported by the Department of Physics and Astronomy, Dartmouth College.
Figure 5: Probability distribution of the ratio of consecutive level spacings for the Hamiltonian (5) in the main text. The blue histogram shows the probability distribution for a system of $N=4$ qubits with the same $M=100$ realizations we use for exact and quantum dynamics. We display the data averaged over $12$ ratios of consequential level spacings without two pairs of states at the edges of the spectrum. The red histogram is done for the system with $N=10$ qubits over $M=1000$ realizations, averaging the data over $\sim 10\%$ out of $2^{10}$ of many-body energy levels. For comparison, the black dashed line shows the GUE predicted level statistics. The gray dotted line displays Poisson (integrable) behavior put for a visual reference. Both ($N=4$, $M=100$) and ($N=10$, $M=1000$) datasets saturate the GUE curve and possess the GUE characteristic value of the mean adjacent gap ratio $\langle r\rangle_{\rm GUE}\approx 0.6$.
Figure 6: Equilibration of the multi-qubit system in the initial state $|\psi_{\rm in}\rangle=|\downarrow\downarrow\ldots\downarrow\rangle$ after the quench. We plot the expectation value of $H_{0}$ as a function of time for individual realizations of random couplings (gray curves), averaged over $M=100$ realizations (solid blue curve), averaged over $M=10$ realizations (dash-dotted green curve), and averaged over $M=20$ realizations (dashed red curve) for different system sizes: a) for $N=4$ qubits, b) for $N=6$ qubits, and c) $N=8$ qubits. The vertical black line marks the equilibration timescale at $t=2.5/J$.
Appendix A Choice of thermalization protocol
The choice of the interaction is essential for our protocol so that there would be no constraints preventing thermalization after the quench. We began with a “constrained” non-interacting Hamiltonian $H_{0}$ that can not be the sole source of thermalization dynamics. In our case, thermalization occurs due to the quenched random interaction ${\cal V}$ (with zero mean strength) between the qubits. Owing to the zero mean strength interaction, adding the latter to the initial Hamiltonian conserves only the total energy equal to the initial energy of the system on average. The inter-qubit interaction $\cal V$ maps onto the random free fermions, also known as the SYK2 model, an explicitly non-interacting theory. However, the combined Hamiltonian $H=H_{0}+{\cal V}$ is non-integrable, as may be seen from its level statistics.
The notion of energy level statistics allows characterizing integrability, localization, and chaos in quantum many-body systems. Chaotic behavior, closely related to thermalization [6], manifests in statistical properties of a system’s many-body energy levels, falling in the Wigner-Dyson statistics described by a Gaussian orthogonal (GOE), unitary (GUE), or symplectic (GSE) ensemble of random matrices [52]. On the contrary, integrability and localization are associated with level statistics following a Poisson distribution [53]. Accordingly, changing disorder in a system may lead to a localization transition from Wigner-Dyson to Poisson level statistics [54, 55, 56, 57, 58, 59].
To quantify the level statistics in our model, we use the probability distribution for a ratio $0\leq r_{l}\leq 1$ determined from two consecutive level spacings in a given part of a many-body spectrum [54]: $r_{l}={\rm min}\{\delta_{l+1},\delta_{l}\}/{\rm max}\{\delta_{l+1},\delta_{l}\}$, where $\delta_{l}=\tilde{\varepsilon}_{l+1}-\tilde{\varepsilon}_{l}$. Note that here $\tilde{\varepsilon}_{l}$ are the eigenvalues of $H$. For the adjacent gap ratio $r_{l}$ following Wigner-Dyson statistics, the probability distribution is $P(r)=2(1+r)^{\beta}/(1+r+r^{2})^{1+\frac{3\beta}{2}}/Z_{\beta}$ with $Z_{\beta}$ being a normalization constant and $\beta$ is a symmetry index (do not confuse with an inverse temperature) equal to $1,2$, and $4$ for GOE, GUE, and GSE, predicting mean values of the gap ratio $\langle r\rangle_{\rm GOE}\approx 0.54$, $\langle r\rangle_{\rm GUE}\approx 0.6$, and $\langle r\rangle_{\rm GSE}\approx 0.68$ [56]. In the integrable case, for the Poisson spectrum, the distribution of the gap ratio is $P(r)=2/(1+r)^{2}$ with $\langle r\rangle_{\rm Poisson}\approx 0.39$ [54]. Following earlier studies [54, 55, 56, 57, 58, 59], we consider our total Hamiltonian $H=H_{0}+{\cal V}$ in the regime when the system is large enough ($N=10$ with $2^{N}$ many-body energy levels) and plot the histogram for the gap ratio distribution for $M=1000$ disorder realizations averaged over $\sim 10\%$ of levels in Fig. 6. The distribution function is in complete agreement with GUE prediction and reproduces the mean adjacent gap ratio $\langle r\rangle_{\rm GUE}\approx 0.6$. Reducing the system’s size and the number of realizations to what we use in our experiment and choosing the parameters accordingly, we observe that the level statistics of the model remain well-described by GUE in Fig. 6, highlighting an absence of integrability or localization.
Having established that the model is chaotic, i.e., described by a random matrix theory, one expects the observables evolved with the Hamiltonian $H$ to equilibrate and thermalize. However, one realization of random couplings may not be enough for equilibration. In Fig. 6, we show the after-the-quench equilibration process for $\langle H_{0}(t)\rangle$ for different system sizes ($N=4,6,8$), averaging the expectation values with several numbers of realizations $M=10,20,100$. Increasing the number of realizations leads to less finite-size oscillations in the average observable in the equilibrium regime. The larger the system, the fewer disorder realizations are needed to suppress the remaining oscillations in the observable. In turn, the time the system evolves to achieve average stationary observables $t_{\rm eq}\sim{\cal O}(1)/J$ does not depend on the system’s size, being reminiscent of a characteristic timescale of irreversible dynamics $T_{3}$ in open quantum systems, solely defined by an inverse interaction amplitude, bypassing which the system relaxes toward thermal equilibrium [51]. This resemblance with open quantum systems is well-aligned with a recent finding that the average time evolution by a random interacting Hamiltonian is equivalent to dissipative dynamics [19]. A complementary observation has been made in the recent work on deep thermalization at infinite temperature tailored for analog quantum simulators, noting that adding a random potential in a simulated Hamiltonian reduces the required amount of the bath qubits [49].
Equilibration in the system results in thermal observables. In particular, we use the probabilities to occupy the many-body energy eigenstates of the initial Hamiltonian $H_{0}$. After the quench, the average occupation probabilities relax to a finite temperature Gibbs distribution. This protocol allows dynamically obtaining equilibrium observables at finite temperatures for a given system on a digital quantum processor without simulating a thermal bath.
Figure 7: Quantum circuit ansatz used for the variational recompilation of the circuit. A 4-qubit unitary $V$ is implemented as a parametrized $m$-layer quantum circuit of alternating odd and even layers. Each universal single-qubit gate $U_{3}$ is parametrized by $3$ distinct parameters, for a total of $n_{p}=12m$ parameters for $V$. Using our recompilation algorithm, we optimize the circuit over the set of $n_{p}$ parameters to minimize the trace distance between $V$ and the target unitary $U(t)=e^{-iHt}$. In our experiments, we use recompiled circuits with $m=20$ layers for a total of $60$ CNOT gates.
Appendix B Variational circuit recompilation
On IBMQ devices, errors largely originate from the application of noisy CNOT gates associated with an error rate of $\sim 1\%$. A naive compilation of the unitary operator $U(t)$ into a quantum circuit using, e.g., Qiskit’s transpile function [60] on a linear qubit layout yields a quantum circuit with as much as $238$ CNOT gates, which would result in a global error rate close to $100\%$. This number of entangling gates originates from the all-to-all interaction (6) essential for our thermalization mechanism. To minimize the number of noisy CNOT gates required for the simulation, we take an alternative approach to recompile our quantum circuits, adapting a variational approach similar to the one presented in Refs. [40, 41], which has already shown promising results on NISQ devices [61, 62, 63]. The general idea of the method is to consider an ansatz quantum circuit with a given CNOT gate structure and parametrized single-qubit gates and optimize these parameters until the circuit implements a suitable approximation of the target unitary $U(t)$. For few-qubits circuits, the optimization is done on a classical computer using a tensor network decomposition of both the ansatz and target circuit.
To perform the recompilation we first generate a parametrized ansatz circuit made of $m$ alternating layers of CNOT and universal $U_{3}$ single-qubit gates on every qubit as shown in Fig. 7. The universal single qubit gate $U_{3}$ gate is characterized by $3$ independent parameters, yielding $3\times 4=12$ independent parameters per layer. Thus, for $N=4$ qubits, the ansatz circuit of $m$ layers produces a 4-qubit unitary $V$, a function of $n_{p}=12m$ parameters. Next, we use the Quimb python library [64] to decompose both $V$ and our target unitary $U(t)$ into rank-2 tensor networks and compute the cost function $C(U,V)=1-\operatorname{Tr}\{V^{\dagger}U\}$. Finally, we perform the optimization for $C$ using the limited-memory Broyden-Fletcher-Goldfarb-Shanno algorithm with box constraints (L-BFGS) [65, 66] in combination with a basin-hopping method [67, 68, 69]. This combination is particularly suitable for optimization problems with many parameters and many local minima. Using this method, we recompile our evolution operator $U(t)$ for each realization of the quench protocol into quantum circuits structured as in Fig. 7 with $m=20$ layers i.e. $60$ CNOT gates with an average fidelity of the order of $1-C(U,V)=99.8\%$.
The optimization runs on a regular laptop in a few minutes. The advantages of this optimization method are a drastic reduction in the number of required CNOT gates by $178$ and no need for a Trotter approximation, which allows us to run the experiment up to arbitrary long times with good precision, crucial to studying thermalization.
Appendix C Error mitigation
Running a time evolution of a quantum-mechanical system on noisy NISQ devices, like IBM quantum computers, is susceptible to noise, generating output errors due to multiple sources. On IBMQ devices, it is known that errors stem principally from imperfect measurements, and from the application of imperfect two-qubit entangling gates, in our case, CNOT gates.
Measurement errors.— To mitigate measurement errors, we execute every circuit with a sampling of $1000$ shots and correct the outcome using the readout error correction technique that involves matrix inversion through the iterative Bayesian unfolding method [70].
Randomized compiling.— To mitigate errors arising from the two-qubit gates application, we aim to simplify the noise structure, which allows for the implementation of efficient error mitigation techniques. To do so, we follow the randomized compiling (RC) approach [71, 72, 73] which enables one to transform any coherent noise affecting CNOT gates into an effective incoherent Pauli noise channel [74]. RC entails generating alternative versions of the quantum circuit of interest, in which every CNOT gate is dressed with Pauli gates acting on individual qubits and drawn randomly from a set of Pauli matrices. These Pauli matrices are drawn in such a way that the dressed CNOT gates are still logically equivalent to an isolated CNOT gate, so that every alternative version of the circuit still implements the same unitary operation. The measured expectation values stemming from each randomly compiled circuit are then averaged, yielding the expectation value that one would have obtained if all CNOT gates were only affected by a purely incoherent noise channel.
In addition to the qubits on which the CNOT gate is applied, a noisy CNOT gate may also lead to coherent errors on neighboring qubits. In our case, this can be formalized by extending the noise channel associated with a given CNOT gate to all qubits in the register. This noise, known as spillover crosstalk noise [75], can also be made purely incoherent by extending the RC method to neighboring qubits. In addition, since a priori no logical operation is being applied on these other qubits, the RC procedure can be improved by applying extra random $\pi/2$ rotations on top of the Pauli matrices, as prescribed in [42]. One can show that by doing this, the effective incoherent noise channel obtained after averaging over crosstalk RC circuits is a combination of a local Pauli noise channel for the active qubits of the CNOT gates, a local depolarizing noise channel acting on the neighboring qubits, and a global depolarizing noise channel acting uniformly on all qubits. In our quantum simulation, we run $100$ crosstalk RC circuits for every realization of the quench protocol, and average the read-out observables, effectively realizing the noise channel described above. We have checked for this number of circuits that the crosstalk RC averages has clearly converged to its infinite sampling value. Using our knowledge of the structure of the effective incoherent noise channel acting on our qubits, we now implement an error mitigation strategy.
Removing global depolarizing noise.— After performing crosstalk RC, we know that at least a part of the effective noise acting on our circuit takes the form of an $N$-qubit global depolarizing noise channel. Due to the simplicity of this noise channel, its effects can be mitigated efficiently with a simple fit, as we show here. For a single CNOT gate, the $N$-qubit global depolarizing noise channel reads [37]:
$\mathcal{E}_{N}(\rho)=f\rho+(1-f)\frac{\mathbb{1}}{2^{N}},$ (12)
where $\rho$ the density matrix, $f$ is the fidelity of a CNOT gate and $\mathbb{1}$ the identity matrix of size $2^{N}\!$. The rightmost term, which models the output of a global depolarizing error occurring, is the maximally mixed state with $N$ qubits. If such an error occurs on any CNOT gate during the execution of the circuit, then the qubits will remain in this state, as it is not affected by the action of subsequent unitary operations. Thus, it is straightforward to generalize Eq. (12) to the case of a circuit with $N_{\text{CNOT}}$ CNOT gates by replacing the fidelity of a single CNOT gate $f$ with a circuit fidelity $f^{N_{\rm CNOT}}$. Further averaging the quantum simulation data for occupation probabilities over $100$ realizations of the quench protocol, we interpret the noisy quantum computer’s output within the model (12), leading to the expression (11) for noisy average occupations in the main text. Then, we leverage the errors due to the local Pauli and depolarizing noise channels using zero-noise extrapolation (ZNE).